Hello,

I need help solving this problem:

Two buildings are separated by a three-foot alleyway. Fatima wants to use a 16-foot ladder to reach a window in the wall of one of these buildings. If she places the foot of the ladder against the base of the other building, how far up the wall will the top of the ladder reach?

Thank you

Answer by stanbon(48535)

Two buildings are separated by a three-foot alleyway. Fatima wants to use a 16-foot ladder to reach a window in the wall of one of these buildings. If she places the foot of the ladder against the base of the other building, how far up the wall will the top of the ladder reach?
---------------------
The ladder is the hypotenuse of a right triangle whose base is 3'
Find the 3rd side using Pythagoras.
-------------------
16^2 = 3^2 + x^2
x^2 = 256 - 9
x = sqrt247
x = 15.716 ft
=====================
Cheers,
Stan H.